1. A point charge q1 is held stationary at the origin. A second charge q2 is placed at point a, and the electric potential energy of the pair of the charges is + 5.4 × 10 raise to -8 J. When the second charge is moved to point b, the electric force does on the charge does -1.9 × 10 raised to -8 J of work. What is the electric potential energy of the pair of charges when the second charge is at point b.

2. How much work is needed to assemble an atomic nucleus containing three protons ( such as Be ) if we model it as an equilateral triangle of side 2.00 × 10 raised to -15 m.
1
Expert's answer
2019-02-05T12:44:07-0500

1. U=U_0-A=5.4∙〖10〗^(-8)+1.9∙〖10〗^(-8)=7.3∙〖10〗^(-8) J

2. A=3∙q^2/(4πε_0 r)=3∙9∙〖10〗^9∙(1.6∙〖10〗^(-19) )^2/(2∙〖10〗^(-15) )=3.456∙〖10〗^(-13) J